OpenStudy (anonymous):

How would I go about solving this? (will give medal & fan) (posting question in comments)

OpenStudy (anonymous):

\[\sin(\frac{ 3\pi }{ 4 })\]

OpenStudy (anonymous):

I know that the thing in the parenthesis is a radian or whatever it's called (the unit aside from degrees used to measure angles)... but I don't exactly know how to find the sine of it...

OpenStudy (anonymous):

1 minute

OpenStudy (anonymous):

can you show me your work to explain how you got that please?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Because 1 minute isn't an option for me...

OpenStudy (anonymous):

\[1/\sqrt{2}\]

OpenStudy (anonymous):

no that wasnt an anser i just said like wait 1 mainute:]

OpenStudy (anonymous):

My options are:\[-\frac{ \sqrt{2} }{ 2 }\]\[\frac{ \sqrt{2} }{ 2 }\]\[\frac{ \sqrt{3} }{ 2 }\]\[\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

yes then it would be \[\sqrt{2}/2\]

OpenStudy (anonymous):

because that is an alternative form of the answer

OpenStudy (anonymous):

oh okay but how did you get that? i just want to know so i know for future reference lol

OpenStudy (anonymous):

sorry dont know i just looked it up on WolframAlpha :/ good luck.

OpenStudy (anonymous):

oh lol that's fine, thanks anyway!